# SmartCooper: Vehicular Collaborative Perception with Adaptive Fusion and Judger Mechanism

### 02 Feb 2024 v2

@@ III-C Overall Architecture
(1) The first stage is network optimization: The network optimization process involves collecting parameters from the actual communication environment to compute the optimized compression ratio $\alpha^{*}$ for both the Adaptive Encoder and Adaptive Decoder. In the case of the Adaptive Encoder and Adaptive Decoder, we have directly implemented the Modulated Autoencoder[25]. This framework allows for a trade-off between bitrate and distortion by utilizing trade-off parameters. These trade-off parameters correspond one-to-one with compression ratios. Therefore, by adjusting the compression ratio, we can influence the trade-off parameters.
@@
(3) The last stage is fusion: We adapt the Fusion Net, i.e. CoBEVT [24], which is designed for the fusion of vehicular camera data. It excels merging data from various vehicles and produces a BEV output.
@@
0:  Input ego CAV’s image sequence: $I_{e}$. CAVs’ image sequence set: $\{I_{i}\}$. Adaptive encoders: $E_{A}$. Adaptive decoders: $D_{A}$. The channel parameters: $N$, $W_{T}$, $P_{T}$, $H$, $\rho_{n}$. The volume of data transmitted: $V$. The vehicle coordinates: $\{C_{0},...,C_{N-1}\}$.
@@
2:     Calculate the adaptive compression ratio ${\alpha^{*}}_{i}$ by Eq. (2);
@@ IV-B Baselines
No Fusion: This scheme does not use any collaborative perception, in which, ego CAV only uses the information it perceives and does not receive data from other CAVs.
@@ IV-C Settings
Our experimental setup complies with 3GPP standards[1]. The basic settings are shown as follows. A vehicle has sensing coverage of up to 70 meters and speeds within the range of 0-50 km/h. The default value of the number of cooperative vehicles is 3, and the upper bound of the number of sub-channels $N$ is 3. Most scenarios in the OPV2V dataset[23] involve collaboration among three or four CAVs. Therefore, the filtering threshold for the judger mechanism is set as follows: In a scenario with three CAVs, the threshold will eliminate data from CAVs whose normalized $score$ is below 0.4 and whose distance from the Ego CAV exceeds $0.5d_{max}$. In a scenario with four CAVs, the threshold will remove data from CAVs whose normalized $score$ falls below 0.3 and whose distance from the Ego CAV is greater than $0.5d_{max}$. Other parameters are set as follows. The default value of the total bandwidth $W_{T}$ is set to 200 MHz, and the default value of the transmission power $P_{T}$ of each vehicle is set to 8 mW. In addition, the road is a bilateral highway with three lanes for vehicles in each direction and with all vehicles uniformly distributed on the road.

### 04 Mar 2024 v3

@@ III-C Overall Architecture
(1) The first stage is network optimization: The network optimization process involves collecting parameters from the actual communication environment to compute the optimized compression ratio $\alpha^{*}$ for both the Adaptive Encoder and Adaptive Decoder. In the case of the Adaptive Encoder and Adaptive Decoder, we have directly implemented the Modulated Autoencoder[29]. This framework allows for a trade-off between bitrate and distortion by utilizing trade-off parameters. These trade-off parameters correspond one-to-one to compression ratios. Therefore, by adjusting the compression ratio, we can adapt the trade-off parameters.
@@
(3) The last stage is fusion: We adapt the Fusion Net, i.e., CoBEVT [28], which is designed for the fusion of vehicular camera data. It conducts merging data from various vehicles and produces a BEV output.
@@
0:  Input ego CAV’s image sequence: $I_{e}$. CAVs’ image sequence set: $\{I_{i}\}$. Adaptive encoders: $E_{A}$. Adaptive decoders: $D_{A}$. The channel parameters: $N$, $W$, $P_{T}$, $H$, $\rho_{n}$. The volume of data transmitted: $V$. The vehicle coordinates: $\{C_{0},...,C_{N-1}\}$.
@@
2:     Calculate the adaptive compression ratio $\alpha_{i}^{*}$ by Eq. (2);
@@ IV-B Baselines
No Fusion: This scheme does not use any collaborative perception, in which ego CAV only uses the information it perceives and does not receive data from other CAVs.
@@ IV-C Settings
Our experimental setup complies with 3GPP standards[1]. The basic settings are shown as follows. A vehicle has sensing coverage of up to 70 meters and speeds within the range of 0-50 km/h. The default value of the number of cooperative vehicles is 3, and the upper bound of the number of sub-channels $N$ is 3. Most scenarios in the OPV2V dataset[27] involve collaboration among three or four CAVs. Therefore, the filtering threshold for the judger mechanism is set as follows: In a scenario with three CAVs, the threshold will eliminate data from CAVs whose normalized $score$ is below 0.4 and whose distance from the Ego CAV exceeds $0.5d_{max}$. In a scenario with four CAVs, the threshold will remove data from CAVs whose normalized $score$ falls below 0.3 and whose distance from the Ego CAV is greater than $0.5d_{max}$. Other parameters are set as follows. The default value of the total bandwidth $W$ is set to 200 MHz, and the default value of the transmission power $P_{T}$ of each vehicle is set to 8 mW. In addition, the road is a bilateral highway with three lanes for vehicles in each direction and with all vehicles uniformly distributed on the road.
@@ V Conclusions
Acknowledgment
This work was supported in part by the Hong Kong SAR Government under the Global STEM Professorship and Research Talent Hub and the Hong Kong Jockey Club under JC STEM Lab of Smart City. The work of X. Chen was supported in part by HKU IDS Research Seed Fund under Grant IDS-RSF2023-0012 and HKU-SCF FinTech Academy R&D Funding.